OpenStudy (anonymous):

You receive two packages of black/white sand mix. Package A is one-third black sand, and two thirds white sand. Package B is nine-tenths black sand, and one tenth white sand. How much of each package is needed to create a mixture that is sixty percent black and forty percent white? I'l like to answer this question myself, but need help in determining how to set it up... any helpers?? please??

OpenStudy (anonymous):

U mean the minimum number of packages

OpenStudy (anonymous):

x*1/3+y*9/10 = 60/100.....(i) x*2/3+y*1/10=40/100...(i)

OpenStudy (anonymous):

SOLVE BOTH THE EQUATIONS

OpenStudy (anonymous):

did u get what I did?

OpenStudy (anonymous):

Why is someone sending me sand?

OpenStudy (anonymous):

Let b be the black sand mix and w be the white send mix then (b/3)+(2w/3)=60% (9b/10)+(w/10)=40% since they are systems of linear equestions with two variables then use simaltanous equations methods. Doyou wantme to carry on?